OpenStudy (anonymous):

a pendulum bob of mass m/2 carrying a charge 2q is at rest with its string making an angle theta with the vertical in uniform horizontal electric field E. the tension in the string is?????

OpenStudy (anonymous):

horzontally, \[T*\sin \theta = 2*q*E\] vertically, \[T*\cos \theta = 1/2*mg\] squaring and adding both equations, \[T^2 * ( (\sin \theta ) ^2 + (\cos \theta)^2 ) = (4*q^2 * E^2) + (1/4 * (mg)^2 )\] hence, \[T^2 = (4*q^2 * E^2) + (1/4 * (mg)^2 )\]

OpenStudy (anonymous):

but ans for the question is 2qE/sin (theta)

OpenStudy (anonymous):

if you note the first line of my answer, you would see that I write the same thing. If you just want the answer in theta, then thats it, no more to do. but if you want to eliminate theta and get the answer in terms of m, E, q etc then you would have to go furthur.

OpenStudy (anonymous):

thank you v v much...

OpenStudy (anonymous):

v v welcome :)